# A Thermodynamically Consistent Free Boundary Model for Two-Phase Flows in an Evolving Domain with Moving Contact Lines, Variable Contact Angles

### v1

@@ 3.3.2 Computations on the boundary
$0&\geq\textnormal{div}_{\Gamma}\big{[}\mathbf{K}_{e}-\theta\mathbf{K}_{\psi}+\partial_{\nabla_{\Gamma}\psi}g\,\partial_{t}^{\circ}\psi-\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}-q{\mathbf{u}}\big{]}\\ &\quad+\big{[}\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big{]}\partial_{t}^{\circ}\phi+\big{[}\partial_{\psi}g-\textnormal{div}_{\Gamma}\big{(}\partial_{\nabla_{\Gamma}\psi}g\big{)}-\theta\big{]}\partial_{t}^{\circ}\psi\\ &\quad+\big{[}\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big{]}_{\tau}\cdot{\mathbf{u}}_{\tau}\\ &\quad+\big{[}\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})\big{]}({\mathbf{u}}\cdot{\mathbf{n}})\\ &\quad+H\mathbf{K}_{e}\cdot{\mathbf{n}}+\nabla_{\Gamma}\theta\,\mathbf{K}_{\psi}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}\end{aligned}$ (3.73)
@@
$0&\geq\big{[}\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big{]}\partial_{t}^{\circ}\phi+\big{[}\partial_{\psi}g-\textnormal{div}_{\Gamma}\big{(}\partial_{\nabla_{\Gamma}\psi}g\big{)}-\theta\big{]}\partial_{t}^{\circ}\psi\\ &\quad+\big{[}\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big{]}_{\tau}\cdot{\mathbf{u}}_{\tau}\\ &\quad+\big{[}\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})+Hq\big{]}({\mathbf{u}}\cdot{\mathbf{n}})\\ &\quad-m_{\Gamma}(\psi)\left|\nabla_{\Gamma}\theta\right|^{2}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}\end{aligned}$ (3.77)
@@
$\big{[}\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big{]}_{\tau}$ $=-\gamma_{\tau}{\mathbf{u}}_{\tau}$ on $\quad\text{on }S_{T},$ (3.79)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})+Hq$ $=-\gamma_{\mathbf{n}}({\mathbf{u}}\cdot{\mathbf{n}})$ on $\quad\text{on }S_{T},$ (3.80)
@@
$0&\geq\big{[}\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big{]}\partial_{t}^{\circ}\phi+\big{[}\partial_{\psi}g-\textnormal{div}_{\Gamma}\big{(}\partial_{\nabla_{\Gamma}\psi}g\big{)}-\theta\big{]}\partial_{t}^{\circ}\psi\\ &\quad-\gamma_{\tau}\left|{\mathbf{u}}_{\tau}\right|^{2}-\gamma_{\mathbf{n}}({\mathbf{u}}\cdot{\mathbf{n}})^{2}-m_{\Gamma}(\psi)\left|\nabla_{\Gamma}\theta\right|^{2}\\ &\quad-(\beta\theta-\mu)\,m_{\Omega}(\phi)\nabla\mu\cdot{\mathbf{n}}\end{aligned}$ (3.81)
@@
    $K\varepsilon\partial_{\mathbf{n}}\phi$ $=\alpha\psi-\phi$ on $\quad\quad\text{on }S_{T}\text{ if $K\in(0,\infty)$},$ (3.83b)
    $\partial_{\mathbf{n}}\phi$ $=0$ on $\quad\quad\text{on }S_{T}\text{ if $K=\infty$},$ (3.83c)
@@
    $\partial_{\phi}g$ $=-\mathds{h}(K)(\alpha\psi-\phi)=-\varepsilon\nabla\phi\cdot{\mathbf{n}}$ on $\quad\text{on }S_{T},$
    $\partial_{\psi}g$ $=\delta^{-1}G^{\prime}(\psi)+\alpha\mathds{h}(K)(\alpha\psi-\phi)=\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\nabla\phi\cdot{\mathbf{n}}$ on $\quad\text{on }S_{T}.$
@@ Acknowledgement
Both authors were partially supported by the “RTG 2339” IntComSin funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 321821685. PK was additionally supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 52469428. YL was partially supported by startup funding from Nanjing Normal University, the Natural Science Foundation of Jiangsu Province (Grant No. BK20240572), and the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (Grant No. 24KJB110020). The support is gratefully acknowledged.

### v2

@@ 3.3.2 Computations on the boundary
$\begin{aligned} 0&\geq\textnormal{div}_{\Gamma}\big[\mathbf{K}_{e}-\theta\mathbf{K}_{\psi}+\partial_{\nabla_{\Gamma}\psi}g\,\partial_{t}^{\circ}\psi-\delta(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi){\mathbf{u}}-q{\mathbf{u}}\big]\\ &\quad+\big[\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big]\partial_{t}^{\circ}\phi+\big[\partial_{\psi}g-\textnormal{div}_{\Gamma}\big(\partial_{\nabla_{\Gamma}\psi}g\big)-\theta\big]\partial_{t}^{\circ}\psi\\ &\quad+\big[\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big]_{\tau}\cdot{\mathbf{u}}_{\tau}\\ &\quad+\big[\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})\big]({\mathbf{u}}\cdot{\mathbf{n}})\\ &\quad+H\mathbf{K}_{e}\cdot{\mathbf{n}}+\nabla_{\Gamma}\theta\,\mathbf{K}_{\psi}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}\end{aligned}$ (3.73)
@@
$\begin{aligned} 0&\geq\big[\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big]\partial_{t}^{\circ}\phi+\big[\partial_{\psi}g-\textnormal{div}_{\Gamma}\big(\partial_{\nabla_{\Gamma}\psi}g\big)-\theta\big]\partial_{t}^{\circ}\psi\\ &\quad+\big[\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big]_{\tau}\cdot{\mathbf{u}}_{\tau}\\ &\quad+\big[\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})+Hq\big]({\mathbf{u}}\cdot{\mathbf{n}})\\ &\quad-m_{\Gamma}(\psi)\left|\nabla_{\Gamma}\theta\right|^{2}+\theta\mathbf{J}_{\psi}\cdot{\mathbf{n}}-\mu\mathbf{J}_{\phi}\cdot{\mathbf{n}}\end{aligned}$ (3.77)
@@
$\big[\mathbf{T}{\mathbf{n}}+\nabla_{\Gamma}q-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}}){\mathbf{u}}+\delta\,\textnormal{div}_{\Gamma}(\nabla_{\Gamma}\psi\otimes\nabla_{\Gamma}\psi)\big]_{\tau}$ $=-\gamma_{\tau}{\mathbf{u}}_{\tau}$ $\quad\text{on }S_{T},$ (3.79)
$\mathbf{T}{\mathbf{n}}\cdot{\mathbf{n}}-\tfrac{1}{2}(\mathbf{J}\cdot{\mathbf{n}})({\mathbf{u}}\cdot{\mathbf{n}})+Hq$ $=-\gamma_{\mathbf{n}}({\mathbf{u}}\cdot{\mathbf{n}})$ $\quad\text{on }S_{T},$ (3.80)
@@
$\begin{aligned} 0&\geq\big[\partial_{\phi}g+\varepsilon\nabla\phi\cdot{\mathbf{n}}\big]\partial_{t}^{\circ}\phi+\big[\partial_{\psi}g-\textnormal{div}_{\Gamma}\big(\partial_{\nabla_{\Gamma}\psi}g\big)-\theta\big]\partial_{t}^{\circ}\psi\\ &\quad-\gamma_{\tau}\left|{\mathbf{u}}_{\tau}\right|^{2}-\gamma_{\mathbf{n}}({\mathbf{u}}\cdot{\mathbf{n}})^{2}-m_{\Gamma}(\psi)\left|\nabla_{\Gamma}\theta\right|^{2}\\ &\quad-(\beta\theta-\mu)\,m_{\Omega}(\phi)\nabla\mu\cdot{\mathbf{n}}\end{aligned}$ (3.81)
@@
    $K\varepsilon\partial_{\mathbf{n}}\phi$ $=\alpha\psi-\phi$ $\quad\quad\text{on }S_{T}\text{ if $K\in(0,\infty)$},$ (3.83b)
    $\partial_{\mathbf{n}}\phi$ $=0$ $\quad\quad\text{on }S_{T}\text{ if $K=\infty$},$ (3.83c)
@@
    $\partial_{\phi}g$ $=-\mathds{h}(K)(\alpha\psi-\phi)=-\varepsilon\nabla\phi\cdot{\mathbf{n}}$ $\quad\text{on }S_{T},$
    $\partial_{\psi}g$ $=\delta^{-1}G^{\prime}(\psi)+\alpha\mathds{h}(K)(\alpha\psi-\phi)=\delta^{-1}G^{\prime}(\psi)+\alpha\varepsilon\nabla\phi\cdot{\mathbf{n}}$ $\quad\text{on }S_{T}.$
@@
4  Conclusion
In this work, we have derived a thermodynamically consistent Navier–Stokes–Cahn–Hilliard model for two-phase flows in an evolving domain with a free boundary. The model incorporates bulk–surface interactions, variable contact angles, and a generalized Navier slip boundary condition, thereby extending previous diffuse-interface formulations to situations involving a domain evolution driven by the fluid motion. A central feature of our framework is that both the bulk and surface phase-fields evolve according to a coupled convective bulk–surface Cahn–Hilliard system, which accounts for material transfer between the bulk and the boundary as well as a variable contact angle.
Starting from local mass and energy balance laws, we systematically derived the governing equations and boundary conditions using the Lagrange multiplier method. The resulting system satisfies the fundamental physical principles of mass conservation, local energy dissipation, and preservation of the bulk volume and surface area of the evolving domain. We further showed that several classical and recently proposed models (cf.[5, 50]) appear as special limiting cases of our formulation. This highlights the generality and flexibility of the present model.
The free boundary setting considered here opens up new perspectives for modeling interfacial phenomena involving deformable substrates, biological membranes, or motile cells. Future work shall address analytical questions such as well-posedness as well as numerical methods capable of approximating the coupled bulk–surface dynamics on evolving geometries. Moreover, applications to concrete physical or biological systems (e.g., droplet spreading on soft surfaces or intracellular flows within moving cells) offer promising directions for further study.
@@ Acknowledgement
Both authors were partially supported by the “RTG 2339” IntComSin funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 321821685. PK was additionally supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 52469428. YL was partially supported by the startup funding from Nanjing Normal University, the Natural Science Foundation of Jiangsu Province (Grant No. BK20240572), the Natural Science Foundation of the Jiangsu Higher Education Institutions of China (Grant No. 24KJB110020), and the China Postdoctoral Science Foundation under Grant Number 2025M773078. The support is gratefully acknowledged.